# Planetarium[Uncaptioned image]: A Rigorous Benchmark for Translating Text to Structured Planning Languages

### v1

@@ Appendix C Implementation of fullySpecify
Appendix D Text-to-PDDL Pair Example
@@ Appendix E License Information
The data for Planetarium is available at https://huggingface.co/datasets/BatsResearch/planetarium and released under a Creative Commons CC-BY-4.0 license. In addition, all code to create the dataset and evaluate models on the benchmark is available at https://github.com/BatsResearch/planetarium and released under a BSD 3 license. The authors are responsible for the content of the dataset.
In our study, we use the Blocksworld and Gripper domains, both of which are sourced from the IPC McDermott [2000]. The specific domain files used in our research are obtained from pddl-generators  Seipp et al. [2022], which is distributed under the GNU General Public License.
Appendix F Additional Experiment Details
Table 2 displays the the hyperparameters used for fine-tuning across all models. In addition, models are loaded using 4-bit NF4 quantization with double quantization to reduce the average memory footprint. We also use bfloat16 compute data type for faster training.
@@ Appendix F Additional Experiment Details
Table 2: Training hyperparameters when fine-tuning all models included in the paper. The hyperparameters are seperated in three parts: supervised fine-tuning related, generation related, and LoRA related.
Appendix G Additional Experiments and Analysis
Figure 5: (Left) Gemma 1.1 IT 2B, fine-tuned on the random train split and evaluated on certain tasks. Most tasks are performed extremely well. (evaluated tasks were seen during training). (Right) Gemma 1.1 IT 2B, fine-tuned on all examples except for the tasks listed above, along with one_room in the initial state and n_room_distributed in the initial state. (tasks not seen during training)
Appendix H Dataset Documentation
To see the most up-to-date documentation, go to the following URL:
https://datanutrition.org/labels/v3/?id=f2b30848-2b93-4f2d-b7e0-926604a02c12
Appendix I Maintenance Plan
The BATS research group at Brown University, will maintain the Planetarium dataset and benchmark with an open GitHub repository and issue submission system, and the dataset hosted on HuggingFace. The maintenance plan includes regular issue tracking, with reviews and categorization of issues, aiming to resolve high-priority issues within a week and scheduling minor updates quarterly. Comprehensive documentation and automated testing will ensure quality and compatibility. Francisco Piedrahita Velez and Max Zuo will lead the maintenance efforts, supported by the BATS research group.

### v2

@@ Appendix B Examples of Algorithm 1
Figure 5 shows examples of how Algorithm 1 checks whether two PDDL problems are equivalent.
@@ Appendix C Implementation of fullySpecify
Floor Tile
In the standard Floor Tile, we know that:
  • A tile cannot be unpainted.
  • All available colors (available-color color) are unchangeable and cannnot be added.
  • Relative positional propositions (right tile1 tile2) and (up tile1 tile2) are unchangeable and cannot be added.
Further, given our simplifications, we have the following characteristics in our domain:
  • Traversal is bidirectional.
  • Painting a tile and other robots do not block path-finding.
  • From the last two facts, robots cannot block themselves off through actions, so the only way the position of a robot is fixed is if:
    1. 1.
      It is explicitly defined in the goal description.
    2. 2.
      The robot is on an isolated tile to begin with.
Applying these rules allows us to fully specify a Floor Tile goal description.
Appendix D Dataset Details
To obtain interesting problems for each domain, we crafted a set of tasks instead of randomly generating instances. Below are descriptions of each domain and their corresponding task configurations.
Figure 6: A visualization of the different tasks in Blocks World. Training set displayed in “magma” color map (purple to yellow). Test set displayed in “viridis” color map (blue to yellow).
Blocks World. This domain involves using a robotic hand to manipulate a set of blocks arranged in various configurations on a table. Blocks World configurations:
  • Stacked: All the blocks are in a single tower.
  • Unstacked: All blocks are separately placed on the table.
  • Holding a Block: All blocks except one are separately placed on the table, the robot holds the last block.
  • Staircase: Towers are in incrementing heights (1, 2, 3, $\ldots$).
  • Equal Towers: Blocks are in towers of equal height.
  • Swap (tied): Goal and initial configurations are tied: given two towers, swap the base blocks, and leave the rest of the structure unchanged.
  • Invert (tied): Goal and initial configurations are tied: Given a set of stacks, flip each stack.
  • Towers: Build towers to match a specified list of heights.
Figure 7: A visualization of the different tasks in Gripper. Training set displayed in “magma” color map (purple to yellow). Test set displayed in “viridis” color map (blue to yellow).
Gripper. This domain features a robot equipped with grippers, tasked with transporting balls from one room to another. Gripper Configurations:
  • Single Room: All balls are in one room.
  • Evenly Distributed: Each room has the same number of balls.
  • Distribute: Each room has a predefined number of balls.
  • Swap Rooms (tied): Goal and initial configurations are tied: swap ball locations between two rooms.
  • Move to Max (goal only): Move all balls to the room that started with the most number of balls.
  • Move to Min (goal only): Move all balls to the room that started with the fewest balls.
  • Pickup (goal only): Pick up all the balls
  • Drop and Pickup (goal only): Drop all the balls in the first room, and pick up all the balls that started in rooms.
  • Juggle (tied): Goal and initial configurations are tied: given that the robot is holding some balls, shuffle which arm holds which ball.
Figure 8: A visualization of the different tasks in Floor Tile. Training set displayed in “magma” color map (purple to yellow). Test set displayed in “viridis” color map (blue to yellow).
Floor Tile. This domain features a group of robots tasked with painting some tiles. Floor Tile configurations:
  • Grid (initial only): Tiles are arranged in a grid.
  • Rings (initial only): Tiles are arranged in a grid in concentric rings.
  • Disconnected Rows (tied): Goal and initial configurations are tied: tiles are in rows, but the rows are disconnected from one another. The goal is to paint both ends of each row.
  • Painted Rings (goal only): Tiles are arranged in a grid in concentric rings, and each ring is painted.
  • Paint All (goal only): All tiles are painted with one color.
  • Painted X (goal only): The robot must paint an X across all the tiles.
  • One Tile, One Color (goal only): Each tile is painted a different color.
  • Checkerboard (goal only): The tiles are painted in a checkerboard pattern with two colors.
@@ Appendix E License Information
The data for Planetarium is available on HuggingFace111hf.co/datasets/BatsResearch/planetarium and released under a Creative Commons CC-BY-4.0 license. In addition, all code to create the dataset and evaluate models on the benchmark is available on GitHub222github.com/BatsResearch/planetarium and released under a BSD 3 license. The authors are responsible for the content of the dataset.
In our study, we use the Blocksworld, Gripper, and Floor Tile domains, which are sourced from the IPC McDermott (2000). The specific domain files used in our research are obtained from pddl-generators  Seipp et al. (2022), which is distributed under the GNU General Public License. The domain file for Floor Tile is a modified version of the one found from pddl-generators  Seipp et al. (2022)
GitHub Copilot was used to assist in writing the code for this paper.
Appendix F Fine-tuning
@@
Table 3: Training hyperparameters when fine-tuning all models included in the paper. The hyperparameters are separated into three parts: supervised fine-tuning, generation-related, and LoRA-related.
We fine-tuned the open-weight models using QLoRA with a rank of $16$, adhering to the hyperparameter recommendations for small models provided by the original authors Dettmers et al. (2023). Models were loaded with 4-bit precision and trained over the training set for a single epoch. Fine-tuning for each model used either two NVIDIA GeForce RTX 3090 GPUs or two NVIDIA A6000 GPUs, operating with data parallelization for approximately 15 hours. We truncate the longest $5\%$ of our training dataset due to GPU memory constraints.
Table 3 displays the hyperparameters used for fine-tuning across all models. In addition, models are loaded using 4-bit NF4 quantization with double quantization to reduce the average memory footprint. We also use the bfloat16 compute data type for faster training.
Appendix G Maintenance Plan
The research group that created Planetarium will maintain the Planetarium dataset and benchmark with an open GitHub repository and issue submission system, and the dataset will be hosted on HuggingFace. The maintenance plan includes regular issue tracking, with reviews and categorization of issues, aiming to resolve high-priority issues within a week and scheduling minor updates quarterly. Comprehensive documentation and automated testing will ensure quality and compatibility.
Appendix H Text-to-PDDL Pair Example